# Performance Comparison of Session-based Recommendation Algorithms based on GNNs

### v1

@@ 2.1 Algorithms
Model
Description
SR-GNN (AAAI ’19)
Maybe the first work that uses a GNN model for SBRS. It constructs session graphs and uses a soft attention mechanism to aggregate information among the items [35].
TAGNN (SIGIR ’20)
Uses an attention module attached to a GNN to adaptively learn the users’ interests [39].
GCE-GNN (SIGIR ’20)
Relies on a two-level item embedding integrated with a position-aware attention module to recursively learn the position of items in a session [33].
COTREC (CIKM ’21)
Based on self-supervised learning with co-training to handle data sparsity and to generate self-supervised signals for SBRS [37].
GNRWW (ICDM ’21)
Combines two kinds of embedding techniques to retrieve local and global patterns from the sessions [43].
FLCSP (J. Inf. Sci. ’21)
Combines latent category abstractions with sequential data to obtain accurate recommendations [42].
CM-HGNN (KBS ’22)
Differently to FLCSP, CM-HGNN leverages actual category information for accurate recommendations [38].
MGS (SIGIR ’22)
Another model [15] which leverages category information. Uses a dual refinement mechanism to move the information between item and category graphs.
@@
Algorithm
Description
SR [18]
Learns sequential rules of order two by considering the co-occurrence of items in a session.
STAN [7]
Builds on SKNN [18] and it uses three decay factors to compute the similarity of sessions and the relevance of the items.
VSTAN [20]
Combines the particularities of VSKNN [18] and STAN and uses a sequence-aware item scoring procedure and an Inverse Document Frequency approach to promote less popular items.
SFSKNN [18]
This variant of SKNN [18] also focuses on the recency of items by considering only those items for recommendations that appear in the neighbor sessions at least once after the last item of the current session.
@@ 3.1 Performance measures
Accuracy. Table 4 shows the accuracy results for the different datasets at the common list lengths of 10 and 20. We sort the results by MRR@20, as this was our optimization target. In terms of the MRR, we find that for all datasets one of the simple methods leads to the highest accuracy values. In particular STAN and VSTAN show consistently good results across several datasets. There is however no single ‘winner’ across datasets. For example, SR works particularly well for the larger RSC15 (1/12) dataset. [^8] The effectiveness of SR on some datasets in terms of MRR was observed also in [18]. Generally, we observe that the differences between the winning simple method and the best GNN-based model are sometimes quite small. Nonetheless, it is surprising that several years after the first publication on the effectiveness of nearest-neighbor techniques in 2017 [12], the most recent published models are not consistently outperforming these baselines.
@@ 3.2 Additional Robustness Analyses for GNN-based Models
The outcomes of these experiments are provided in Table 6.111111Note: the MRR@20 values of some GNN models are not the same as in Table 4 because we decreased the values of batch sizes to avoid memory issues, which in turn affects the accuracy of the models. We can observe that some models, depending on the embedding size, can lead to largely different results in terms of the MRR@20. This confirms the importance of carefully tuning this central hyperparameter. As the results show, the differences between the best and worst MRR values can be more than 40% in some cases, simply by choosing a bad value for the embedding size. For some of the models, the difference between the minimum and maximum MRR values is less pronounced, but still in a range where researchers would report a substantial improvement over the state of the art when comparing models. We note that no clear pattern was found across the models that larger embedding sizes are consistently better than smaller ones.
@@
Random seeds. The random seed used in an experiment is certainly not a hyperparameter to tune. Still, its choice can have notable impact on accuracy results ([34, 23]). We therefore conducted experiments to analyze to what extent this is true also for some of the examined GNN-based models. In these experiments, we selected random seed values between 100 to 10,000,000 by using a random function. In the shared repositories of the GNN models, random seed values between 2000 to 3000 can often been found, but there is no general pattern. In our experiments, we deliberately picked a larger range to examine if some more uncommon values (outliers) could have an unexpected effect. Again, we considered the RSC15 (1/64) and the DIGI dataset. We generated 100 random seed values for the RSC15 dataset, and 35 values for the larger DIGI dataset. The hyperparameters of the examined models were again left at their optimum values.
The obtained results for MRR@20 when using different random seeds are shown in Table 7, again sorted by the difference between the minimum and the maximum observed value. Again, we can see that for some models and datasets, the obtained values using the same set of optimized hyperparameters vary strongly. For the RSC15 dataset and the FLCSP method, for example, the obtained accuracy results can be 30% lower than the best value simply because of a unlucky choice of the random seed value. [^12] In this case, the random seed value leading to the worst MRR value was around 7,000,000. All detailed results can be found online. Clearly, the differences are not as extreme for all models and datasets and may be partly due to the large range of values that we explored. The distribution of the accuracy values Figure 1 for the RSC15 (1/64) dataset. The detailed data of the entire experiment can be found in the online material.
@@
Comparing the results of Table 7 and the results obtained when using the random seed values used in the original code shared by the authors (Table 4), we observe that often higher accuracy values can be achieved when exploring a large number of random seed values. For example, the optimal MRR@20 value for the FLCSP method on the RSC15 (1/64) dataset is 0.183. From the first row in Table 7 we can see that the worst MRR value was at 0.159 and the best one at 0.250, which can be seen as a substantial improvement, which is only obtained by changing the random seed. A similar observation can be made in other situations as well, e.g., considering the last row of Table 7, where the best MRR@20 value of GNRWW is at 0.352 for the DIGI dataset, whereas the best value in Table 4 was at 0.324. Note, however that in some cases the best MRR values obtained by exploring many (somewhat unusual) seed values are lower than when using more common choices as can be found in the shared code repositories, which means that these common ranges should be explored as well.

### v2

@@ 2.1 Algorithms
Model Description
SR-GNN (AAAI ’19) Maybe the first work that uses a GNN model for SBRS. It constructs session graphs and uses a soft attention mechanism to aggregate information among the items [35].
TAGNN (SIGIR ’20) Uses an attention module attached to a GNN to adaptively learn the users’ interests [39].
GCE-GNN (SIGIR ’20) Relies on a two-level item embedding integrated with a position-aware attention module to recursively learn the position of items in a session [33].
COTREC (CIKM ’21) Based on self-supervised learning with co-training to handle data sparsity and to generate self-supervised signals for SBRS [37].
GNRWW (ICDM ’21) Combines two kinds of embedding techniques to retrieve local and global patterns from the sessions [43].
FLCSP (J. Inf. Sci. ’21) Combines latent category abstractions with sequential data to obtain accurate recommendations [42].
CM-HGNN (KBS ’22) Differently to FLCSP, CM-HGNN leverages actual category information for accurate recommendations [38].
MGS (SIGIR ’22) Another model [15] which leverages category information. Uses a dual refinement mechanism to move the information between item and category graphs.
@@
Algorithm Description
SR [18] Learns sequential rules of order two by considering the co-occurrence of items in a session.
STAN [7] Builds on SKNN [18] and it uses three decay factors to compute the similarity of sessions and the relevance of the items.
VSTAN [20] Combines the particularities of VSKNN [18] and STAN and uses a sequence-aware item scoring procedure and an Inverse Document Frequency approach to promote less popular items.
SFSKNN [18] This variant of SKNN [18] also focuses on the recency of items by considering only those items for recommendations that appear in the neighbor sessions at least once after the last item of the current session.
@@ 3.1 Performance measures
Accuracy. Table LABEL:tab:accuracyResults shows the accuracy results for the different datasets at the common list lengths of 10 and 20. We sort the results by MRR@20, as this was our optimization target. In terms of the MRR, we find that for all datasets one of the simple methods leads to the highest accuracy values. In particular STAN and VSTAN show consistently good results across several datasets. There is however no single ‘winner’ across datasets. For example, SR works particularly well for the larger RSC15 (1/12) dataset. [^8] The effectiveness of SR on some datasets in terms of MRR was observed also in [18]. Generally, we observe that the differences between the winning simple method and the best GNN-based model are sometimes quite small. Nonetheless, it is surprising that several years after the first publication on the effectiveness of nearest-neighbor techniques in 2017 [12], the most recent published models are not consistently outperforming these baselines.
@@
Table 4: Accuracy results, sorted by MRR@20. (cont.).
An ANOVA analysis revealed that there are significant differences (p$<$0.05) for the obtained MRR@20 values for the DIGI and RETAIL datasets across all examined models. However, the differences between the best-performing models for these datasets is too small to reach statistical significance according to a Tukey post-hoc test. Nonetheless, given these results, we can conclude that the best simple models are performing at least equally well as the neural ones.
@@ 3.2 Additional Robustness Analyses for GNN-based Models
The outcomes of these experiments are provided in Table 6.111111Note: the MRR@20 values of some GNN models are not the same as in Table 4 because we decreased the values of batch sizes to avoid memory issues, which in turn affects the accuracy of the models. We can observe that some models, depending on the embedding size, can lead to largely different results in terms of the MRR@20. This confirms the importance of carefully tuning this central hyperparameter. As the results show, the differences between the best and worst MRR values can be more than 40% in some cases, simply by choosing a bad value for the embedding size. For some of the models, the difference between the minimum and maximum MRR values is less pronounced, but still in a range where researchers would report a substantial improvement over the state of the art when comparing models. We note that no clear pattern was found across the models that larger embedding sizes are consistently better than smaller ones. Furthermore, the effect size ($\eta^{2}$) is measured using the ANOVA test and found that its values for the DIGI and RSC15 datasets are 85.53% and 91.39% respectively, indicating the high variance between the mean MRR values of the GNN models.
@@
Random seeds. The random seed used in an experiment is certainly not a hyperparameter to tune. Still, its choice can have notable impact on accuracy results ([34, 23]). We therefore conducted experiments to analyze to what extent this is true also for some of the examined GNN-based models. In these experiments, we selected random seed values between 100 to 10,000,000 by using a random function. In the shared repositories of the GNN models, random seed values between 2000 to 3000 can often been found, but there is no general pattern. In our experiments, we deliberately picked a larger range to examine if some more uncommon values (outliers) could have an unexpected effect. Again, we considered the RSC15 (1/64) and the DIGI dataset. We generated 100 random seed values for the RSC15 dataset, and 35 values for the larger DIGI dataset. The hyperparameters of the examined models were again left at their optimum values. The obtained results for MRR@20 when using different random seeds are shown in Table 7, again sorted by the difference between the minimum and the maximum observed value. Again, we can see that for some models and datasets, the obtained values using the same set of optimized hyperparameters vary strongly. For the RSC15 dataset and the FLCSP method, for example, the obtained accuracy results can be 30% lower than the best value simply because of a unlucky choice of the random seed value. [^12] In this case, the random seed value leading to the worst MRR value was around 7,000,000. All detailed results can be found online. Clearly, the differences are not as extreme for all models and datasets and may be partly due to the large range of values that we explored. The distribution of the accuracy values Figure 1 for the RSC15 (1/64) dataset. The detailed data of the entire experiment can be found in the online material.
@@
Comparing the results of Table 7 and the results obtained when using the random seed values used in the original code shared by the authors (Table LABEL:tab:accuracyResults), we observe that often higher accuracy values can be achieved when exploring a large number of random seed values. For example, the optimal MRR@20 value for the FLCSP method on the RSC15 (1/64) dataset is 0.183. From the first row in Table 7 we can see that the worst MRR value was at 0.159 and the best one at 0.250, which can be seen as a substantial improvement, which is only obtained by changing the random seed. A similar observation can be made in other situations as well, e.g., considering the last row of Table 7, where the best MRR@20 value of GNRWW is at 0.352 for the DIGI dataset, whereas the best value in Table LABEL:tab:accuracyResults was at 0.324. Note, however that in some cases the best MRR values obtained by exploring many (somewhat unusual) seed values are lower than when using more common choices as can be found in the shared code repositories, which means that these common ranges should be explored as well. Furthermore, $\eta^{2}$ is measured and found to be 97.79% and 89.56% for the DIGI and RSC15 datasets respectively, indicating the high variance between the mean MRR values of the GNN models.